# Dynamic Investment-Driven Insurance Pricing: Equilibrium Analysis and Welfare Implication

### v1

@@ Abstract
This paper develops a dynamic model to analyze the general equilibrium of the insurance market, focusing on the interaction between insurers’ underwriting and investment strategies. Three possible equilibrium outcomes are identified: a positive insurance market, a zero insurance market, and market failure. Our findings reveal why insurers may rationally accept underwriting losses by setting a negative safety loading while relying on investment profits, particularly when there is a negative correlation between insurance gains and financial returns. Additionally, we explore the impact of regulatory frictions, showing that while imposing a cost on investment can enhance social welfare under certain conditions, it may not always be necessary. Therefore, we emphasize the importance of tailoring regulatory interventions to specific market conditions.
Keywords: Equilibrium Insurance Pricing; Negative Safety Loading; Social Welfare; Regulatory Friction
@@ 1 Introduction
In 2023, it is reported that among the 69 non-listed property and casualty (P&C) insurance companies in China, 43 of them (over 60%) experienced underwriting losses. [^1] See https://baijiahao.baidu.com/s?id=1790310867880623315&wfr=spider&for=pc, where the research report was published by Thirteen Actuaries, one of China’s largest insurance data consulting firms. Besides, the median underwriting profit margin across these insurers was -3.39%, while the simple average stood at -5.21%, and the weighted average at -0.59%. These figures collectively reflect that the underwriting operations within the P&C industry were largely unprofitable. This phenomenon stands in stark contrast to traditional textbook theories of insurance economics (see Zweifel et al.,, 2021), which posit that insurance pricing usually consists of both an actuarially fair premium and a loading premium (also known as the safety loading), ensuring that the expected profits of insurers are always non-negative. [^2] Based on the Law of Large Numbers, the actuarially fair premium could cover expected losses, while the loading premium would account for operational costs and profit margins, allowing insurers to generate positive expected returns.
@@
Despite the widespread underwriting losses, most of these P&C insurers (over 90% of the total) reported positive investment returns. Statistics show that the median total investment return for these insurers was 3.0%, with a simple average of 2.6% and a maximum of 6.22%. [^3] Also see https://baijiahao.baidu.com/s?id=1790310867880623315&wfr=spider&for=pc. Different from banks, mutual funds, or even life insurers, the primary role of P&C insurers should be that of risk management institutions. However, it is evident that they are increasingly behaving like other financial intermediaries, heavily reliant on investment returns. This trend is consistent with what has been observed in the U.S. insurance market, where insurers have long been the largest institutional investors in corporate bonds (Koijen and Yogo, 2023b, ). Beyond fixed-income investments, insurers are also deeply involved in equity investments. A notable example is Warren Buffett’s investment strategy through Berkshire Hathaway, where the float generated by insurers has been used to finance large-scale equity investments. This approach has allowed Berkshire Hathaway to become one of the largest shareholders in several major corporations.
@@
In this paper, we propose a dynamic framework to study the general equilibrium of the insurance market. The theoretical results we derive under certain conditions can provide an explanation for the two phenomena mentioned above: insurers may rationally choose to incur underwriting losses while relying on investment profits to sustain their operations in equilibrium. More specifically, unlike the traditional insurance economics and actuarial science literature where the safety loading is treated as an exogenously given and positive constant, our model regards the loading as the price of insurance, determined endogenously by the interaction of market supply and demand. Such a pricing method can be referred to as the equilibrium pricing of insurance. Within this framework, we can further extend the analysis to examine the social welfare implications of the insurance market and its relationship with the financial market.
@@
While these studies all conclude that insurers should realize positive expected profits in policies due to a positive safety loading in equilibrium, we propose that the equilibrium loading can naturally become negative under certain parameter settings when we introduce an external financial market for insurers to invest in and the investment returns are negatively correlated with uncertain underwriting gains. [^4] Negative correlation between investment returns and underwriting gains is equivalent to positive correlation between investment returns and underwriting losses. In terms of expression, we also refer to this scenario as a negative correlation between the two markets or risks in this paper. In such a scenario, insurers may be incentivized to offer negative loading premiums to attract policyholders, as the expected returns from investments outweigh the underwriting losses. By leveraging premiums as investment capital, insurers can maintain profitability despite incurring underwriting deficits. [^5] This property was hinted at in Henriet et al., (2016), which slightly discusses the possibility of negative equilibrium loading as insurers hedge their losses through complementary financial market activities. However, they do not provide a detailed proof or specify the market conditions under which this could occur. Our work extends and refines their discussion. And this property helps explain the phenomenon observed among Chinese P&C insurers.
In our setting, introducing a correlation between risks in insurance and financial markets is both natural and necessary, as financial market performance and insurance outcomes are often intertwined. [^6] For instance, during economic downturns insurers typically experience lower investment returns coupled with higher underwriting losses, such as increased property claims or credit defaults, as seen in the 2008 financial crisis. Conversely, during economic booms, while financial markets perform strongly, insurance losses, particularly in liability and workers’ compensation, may increase due to heightened economic activity, such as construction related accidents. Climate events like hurricanes not only lead to substantial insurance claims but can also disrupt financial markets, creating volatility. Although some studies have examined how the correlation between the two risks affects insurers’ investment behavior (see Cheng et al.,, 2024), few have explored its impact on underwriting and pricing strategies. When the two risks driven by Brownian motions are uncorrelated, the insurance and financial markets function as independent entities, and the insurer’s optimal underwriting strategy remains consistent with the scenario where no investment opportunities exist. It can be proved that if an equilibrium exists in this market, the equilibrium loading will be strictly positive.
@@
Within this framework, we further study the impact of regulatory frictions on market equilibrium and social welfare. In practice, insurers’ investment behavior is constrained by risk-based capital regulation, which limits their risk-shifting incentives (see Koijen and Yogo,, 2022). To capture this, we introduce a quadratic form of regulatory-induced investment costs into the insurer’s capital flow process and similarly solve for the market equilibrium. In a positive insurance market, regulatory costs do not affect underwriting strategy, market equilibrium, or social welfare when the two markets are uncorrelated. However, when the two markets are correlated, regulatory costs would affect these outcomes, and the optimal cost that maximizes social welfare can be derived, which is not necessarily zero. Therefore, while in certain cases appropriate regulation of insurers’ investment behavior can enhance social welfare in the insurance market, in other situations regulation may be unnecessary. A case-by-case analysis is required for policy makers.
We contribute to the growing literature on the financial economics of insurance. While classical insurance market theories have primarily focused on the demand side, recent studies have emphasized the importance of supply-side dynamics, highlighting the role of insurers as financial intermediaries in capital markets (Koijen and Yogo, 2023a, ). [^7] For more related literature, see Ge, (2022), Egan et al., (2022), Ellul et al., (2022), Sen, (2023), Kubitza et al., (2023), Koijen et al., (2024), and others. Financial and product market frictions, combined with statutory reserve regulations, can lead to the shadow cost and result in higher prices (Koijen and Yogo,, 2015, 2022). [^8] As a result, insurers are incentivized to reduce capital requirements through shadow reinsurance mechanisms (Koijen and Yogo,, 2016). Beyond pricing, insurers’ balance sheets significantly impact asset prices, which in turn affect the financing and investment decisions of real sector firms (Kubitza,, 2021). While several empirical studies exist, there is limited work on the theoretical foundations of the financial economics of insurance. Our framework can not only capture the dynamic equilibrium of insurance markets and explain the negative loading phenomenon, but also allow for further analysis of social welfare and regulatory implications.
Knox and Sørensen, (2024) is the closest to this paper, as it also studies insurers’ investment strategies and their impact on insurance pricing. We focus on different aspects of insurers’ investment. While asset prices are endogenously determined by insurers’ liquidity demands in their discrete three-period model, we adopt a classic continuous-time finance model (where asset prices are exogenously given by a Black-Scholes market assumption) to analyze the effects of the correlation between risks in the insurance and financial markets.
@@
We also contribute to the ongoing literature in actuarial science on optimal investment and pricing for insurers. While previous studies have examined insurers’ investment strategies under various optimization objectives (Liu and Yang,, 2004; Yang and Zhang,, 2005), with ambiguity aversion (Zeng et al.,, 2016; Guan and Liang,, 2019), and in connection with optimal reinsurance (Zeng and Li,, 2011; Yi et al.,, 2013), our focus lies in the interaction between investment and underwriting strategies. Regarding insurance pricing, while fewer studies address how premiums should be determined in a competitive market, one approach is to model it as a game between insurers, as explored in Emms and Haberman, (2005), Emms, (2007, 2012), Asimit and Boonen, (2018) and Li et al., (2021). Instead, we adopt a representative insurer model and follow the equilibrium pricing framework of insurance proposed by Henriet et al., (2016) and Feng et al., (2023), though we believe that the game-theoretical problem within this framework merits further research.
@@ 2 General Model and Problem Formulation
Let $\Big{(}\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P}\Big{)}$ be a complete probability space satisfying the usual conditions. All stochastic processes governing the insurance and financial markets are assumed to be well-defined on this probability space.
@@ 2.1 Insurance Market
Consider a competitive insurance market that consists of a continuum of insurers, each offering insurance to individuals exposed to perfectly correlated risks. The idiosyncratic risks of individuals are ignored, as they can be diversified away. The cumulative loss process for a representative insurer, denoted by $L\triangleq\Big{\{}L_{t}:t\geq 0\Big{\}}$, evolves as the following dynamics:
@@
where $l>0$ represents the expected instantaneous loss rate, $\eta>0$ denotes the exposure level to the systematic loss risk, and $W^{I}\triangleq\Big{\{}W^{I}_{t}:t\geq 0\Big{\}}$ is a one-dimensional standard Brownian motion modeling the risk. [^9] In principle, the systematic loss risk may stem from multiple sources and can be modeled as multi-dimensional Brownian motion. For simplicity, and under the assumption of homogeneity in insurers’ risk exposures, we aggregate them into a single dimension for tractability.
@@
Specifically, on the supply side, insurers determine their underwriting amount based on the prevailing insurance price. The underwriting amount at time $t$, denoted by $x_{t}\geq 0$, constitutes the underwriting process $X\triangleq\Big{\{}x_{t}:t\geq 0\Big{\}}$. Then the insurer’s surplus process $M=\{m_{t}:t\geq 0\}$ evolves as the following dynamics:
@@
and the equilibrium price $\theta^{\ast}_{t}$ is solved endogenously. The price process is denoted by $\Theta\triangleq\Big{\{}\theta_{t}:t\geq 0\Big{\}}$, while the demand process is denoted by $D\triangleq\Big{\{}d_{t}=d(\theta_{t}):t\geq 0\Big{\}}$.
@@ 2.2 Financial Market
where $r$ is the risk-free rate, $\mu>r$ represents the expected return on the risky asset, and $\sigma>0$ denotes the volatility of the return. Here, $W^{S}\triangleq\Big{\{}W^{S}_{t}:t\geq 0\Big{\}}$ is a one-dimensional standard Brownian motion, which may be correlated with the insurance loss risk $W^{I}$. The correlation coefficient is given by $\rho\in(-1,1)$, such that $\mathrm{d}W^{I}_{t}\mathrm{d}W^{S}_{t}=\rho\mathrm{d}t$. Here we exclude the possibility of $\rho=\pm 1$ to simplify our analysis. In practice, perfect correlation between insurance and financial risks is almost impossible, as that would imply that one of the two markets would lose its distinct purpose. But we can still analyze the implications of such extreme cases from a theoretical perspective, see Subsection 3.7.
@@
Suppose that the insurer follows an investment strategy $Y\triangleq\Big{\{}y_{t}:t\geq 0\Big{\}}$, putting $y_{t}$ amount in the risky asset at time $t$, and the remaining $m_{t}-y_{t}$ in the risk-free asset. As a result, the insurer’s wealth process $M\triangleq\Big{\{}m_{t}:t\geq 0\Big{\}}$ follows:
$\mathrm{d}m_{t}=\Big{(}x_{t}\theta_{t}l+y_{t}(\mu-r)+m_{t}r\Big{)}\mathrm{d}t+x_{t}\eta\mathrm{d}W^{I}_{t}+y_{t}\sigma\mathrm{d}W^{S}_{t}.$ (2.1)

### v2

# Dynamic Investment-Driven Insurance Pricing and Optimal Regulation

@@ Abstract
This paper analyzes the equilibrium of insurance market in a dynamic setting, focusing on the interaction between insurers’ underwriting and investment strategies. Three possible equilibrium outcomes are identified: a positive insurance market, a zero insurance market, and market failure. Our findings reveal why insurers may rationally accept underwriting losses by setting a negative safety loading while relying on investment profits, particularly when there is a negative correlation between insurance gains and financial returns. Additionally, we explore the impact of regulatory frictions, showing that while imposing a cost on investment can enhance social welfare under certain conditions, it may not always be necessary.
Keywords: Equilibrium insurance pricing; Insurer investment; Social welfare; Insurance regulation
@@ 1 Introduction
The question of how insurers price their products reasonably and how regulators oversee these companies effectively has been a long-standing area of focus for both academia and industry, particularly within the field of insurance economics and actuarial science. Theoretically, insurers would not face solvency issues under the principle of actuarially fair pricing, as insurance pricing usually consists of both an actuarially fair premium and a loading premium (also known as the safety loading), ensuring that the expected profits of insurers are always non-negative. [^1] Based on the Law of Large Numbers, the actuarially fair premium could cover expected losses, while the loading premium would account for operational costs and profit margins, allowing insurers to generate positive expected returns. Solvency issues typically arise only in the event of unforeseen extreme risk losses or interest rate changes, but such fluctuations are inherently short-term. [^2] This is because, after losses occur, insurers can identify the underlying causes and dynamically adjust their pricing to avoid negative expected future profits, for instance, by revising risk models or updating parameter estimations.
However, recent practices reveal that underwriting losses are becoming increasingly common among insurers. For instance, in 2023, it was reported that among the 69 non-listed property and casualty (P&C) insurers in China, 43 of them (over 60%) experienced underwriting losses. [^3] See https://baijiahao.baidu.com/s?id=1790310867880623315&wfr=spider&for=pc, where the research report was published by Thirteen Actuaries, one of China’s largest insurance data consulting firms. It is reported that the median underwriting profit margin across these insurers was -3.39%, while the simple average stood at -5.21%, and the weighted average at -0.59%. Moreover, these losses appear to be persistent rather than short-term. According to a 2022 research report by the China Insurance Security Fund, over half of the P&C insurers in China have experienced underwriting losses for five consecutive years. [^4] The China Insurance Security Fund is a state-owned enterprise responsible for managing the insurance security fund and monitoring risk in the insurance industry. This conclusion is derived from its report China Insurance Industry Risk Assessment Report 2022. Figure 1.1 illustrates the operating performance of a representative mid-sized insurer from 2019 to 2023.555The data is sourced from the company’s annual information disclosure reports, which are publicly available and can be downloaded from https://www.cjbx.com.cn/ndxxplbg/index.jhtml. Its three primary business lines all incurred losses every year during this period. The phenomenon of largely unprofitable underwriting operations within the P&C industry stands in stark contrast to traditional insurance pricing theory. Understanding why and how this occurs is both important and necessary for informing industry practices.
Figure 1.1: Investment Gains and Underwriting Losses of a Representative Insurer
Despite the widespread underwriting losses, in reality, most P&C insurers manage to generate profits through investments. For example, among the aforementioned 69 P&C insurers in China, over 90% reported positive investment returns in 2023.666Also see https://baijiahao.baidu.com/s?id=1790310867880623315&wfr=spider&for=pc, where the research report was published by Thirteen Actuaries, one of China’s largest insurance data consulting firms. Statistics indicate that the median total investment return for these insurers was 3.0%, with a simple average of 2.6% and a maximum of 6.22%. Specifically, the insurer in Figure 1.1 achieved investment gains over the past five years, although the gains were insufficient to fully offset underwriting losses. Moreover, there appears to be a negative relationship between the insurer’s underwriting gains and investment returns. In years with higher aggregate underwriting losses, such as 2021, the insurer recorded its highest investment returns. Conversely, in years with relatively lower aggregate underwriting losses, such as 2019 and 2023, the insurer’s investment returns were also relatively lower. More precisely, it is calculated that the correlation coefficient between the insurer’s investment gains and the underwriting gains of its three main business lines is -0.79 over the past five years.
@@
Different from banks, mutual funds, or even life insurers, the primary role of P&C insurers should be that of risk management institutions. However, it is evident that they are increasingly behaving like other financial intermediaries, heavily reliant on investment returns. This trend is also consistent with what has been observed in the U.S. insurance market, where insurers have long been the largest institutional investors in corporate bonds (Koijen and Yogo, 2023b, ). Beyond fixed-income investments, insurers are also deeply involved in equity investments. A notable example is Warren Buffett’s investment strategy through Berkshire Hathaway, where the float generated by insurers has been used to finance large-scale equity investments. This approach has allowed Berkshire Hathaway to become one of the largest shareholders in several major corporations.
These observations motivate us to investigate the incentives behind insurers’ underwriting and investment decisions. In this paper, we adapt the dynamic framework of Henriet et al., (2016) to analyze the equilibrium outcomes of the insurance market. The theoretical results derived under specific conditions indicate that insurers may rationally choose to incur underwriting losses while relying on investment profits to sustain their operations in a long-term equilibrium, driven by the diversification effect. This property provides a possible explanation for the phenomenon observed among Chinese P&C insurers, as discussed.
@@
More specifically, unlike the traditional approach where the safety loading is treated as an exogenously given and positive constant, our model considers the loading as the price of insurance, endogenously determined by the interaction of market supply and demand. Such a pricing method can be referred to as the equilibrium pricing of insurance. Within this framework, we further explore the welfare implications of the insurance market and its interplay with the financial market. Additionally, recognizing the importance of regulatory oversight, we examine the optimal regulatory policies that could be implemented under varying market conditions to address insurers’ behaviors effectively.
@@
While these studies all conclude that insurers should realize positive expected profits in policies due to a positive safety loading in equilibrium, we propose that the equilibrium loading can naturally become negative under certain parameter settings when we introduce an external financial market for insurers to invest in and the investment returns are negatively correlated with uncertain underwriting gains. [^7] Negative correlation between investment returns and underwriting gains is equivalent to positive correlation between investment returns and underwriting losses. In terms of expression, we also refer to this scenario as a negative correlation between the two markets or risks in this paper. In such a scenario, insurers may be incentivized to offer negative loading premiums to attract policyholders, as the expected returns from investments outweigh the underwriting losses. By leveraging premiums as investment capital, insurers can maintain profitability despite incurring underwriting deficits. [^8] This property was hinted at in Henriet et al., (2016), which slightly discusses the possibility of negative equilibrium loading as insurers hedge their losses through complementary financial market activities. However, they do not provide a detailed proof or specify the market conditions under which this could occur. Our work extends and refines their discussion.
In our setting, introducing a correlation between risks in insurance and financial markets is both natural and necessary, as financial market performance and insurance outcomes are often intertwined. [^9] For instance, during economic downturns insurers typically experience lower investment returns coupled with higher underwriting losses, such as increased property claims or credit defaults, as seen in the 2008 financial crisis. Conversely, during economic booms, while financial markets perform strongly, insurance losses, particularly in liability and workers’ compensation, may increase due to heightened economic activity, such as construction related accidents. Climate events like hurricanes not only lead to substantial insurance claims but can also disrupt financial markets, creating volatility. Although some studies have examined how the correlation between the two risks affects insurers’ investment and reinsurance behavior (see Cheng et al.,, 2024; Wu et al.,, 2025), few have explored its impact on underwriting and pricing strategies. When the two risks driven by Brownian motions are uncorrelated, the insurance and financial markets function as independent entities, and the insurer’s underwriting strategy remains consistent with the scenario where no investment opportunities exist.
@@
Within this framework, we further study the impact of regulatory frictions on market equilibrium and social welfare. In practice, insurers’ investment behavior is constrained by risk-based capital regulation, which limits their risk-shifting incentives (see Koijen and Yogo,, 2022). To capture this, we introduce a quadratic form of regulatory-induced investment costs into the insurer’s capital flow process and similarly solve for the equilibrium. In a positive insurance market, regulatory costs do not affect underwriting strategy, market equilibrium, or social welfare when the two markets are uncorrelated. However, when the two markets are correlated, regulatory costs would affect these outcomes, and the optimal cost that maximizes social welfare can be derived, which is not necessarily zero. Therefore, while in certain cases appropriate regulation of insurers’ investment behavior can enhance social welfare in the insurance market, in other situations regulation may be unnecessary. A case-by-case analysis is required for policy makers.
@@
We contribute to the ongoing literature in actuarial science on optimal investment and pricing for insurers. While previous studies have examined insurers’ investment strategies under various optimization objectives (Liu and Yang,, 2004; Yang and Zhang,, 2005), with ambiguity aversion (Zeng et al.,, 2016; Guan and Liang,, 2019), and in connection with optimal reinsurance (Zeng and Li,, 2011; Yi et al.,, 2013), our focus lies in the interaction between investment and underwriting strategies. Regarding insurance pricing, while fewer studies address how premiums should be determined in a competitive market, one approach is to model it as a game between insurers, as explored in Emms and Haberman, (2005), Emms, (2007, 2012), Asimit and Boonen, (2018) and Li et al., (2021). Instead, we adopt a representative insurer model and follow the equilibrium pricing framework of insurance proposed by Henriet et al., (2016) and Feng et al., (2023), though we believe that the game-theoretical problem within this framework merits further research.
We also contribute to the growing literature on the financial economics of insurance. While classical insurance market theories have primarily focused on the demand side, recent studies have emphasized the importance of supply-side dynamics, highlighting the role of insurers as financial intermediaries in capital markets (Koijen and Yogo, 2023a, ). [^10] For more related literature, see Ge, (2022), Egan et al., (2022), Ellul et al., (2022), Sen, (2023), Kubitza, (2021), Kubitza et al., (2023), Koijen et al., (2024), Knox and Sørensen, (2024), and others. Financial and product market frictions, combined with statutory reserve regulations, can lead to the shadow cost and result in higher prices (Koijen and Yogo,, 2015, 2022). While several empirical studies exist, there is limited work on the theoretical foundations of the financial economics of insurance. Our framework not only capture the dynamic equilibrium of insurance markets and explain the negative loading phenomenon, but also allow for further analysis of social welfare and regulatory implications.
We must acknowledge the limitations of this paper. Theoretically, the prices of financial assets follow the classical Black-Scholes market assumption and are given exogenously in our model. It would be interesting to explore how insurers influence the prices of financial assets in a general equilibrium model. Empirically, due to data limitations, we do not provide a detailed investigation into the relationship between the profitability of insurers’ business and investment activities. As the main focus of this paper is to derive theoretical insights, we just cite statistics from research reports and present data from a representative insurer to support our argument, which lacks rigorous empirical validation. Addressing these limitations could be a promising direction for future studies.
@@ 2 General Model and Problem Formulation
Let $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\}_{t\geq 0},\mathbb{P})$ be a complete probability space satisfying the usual conditions. All stochastic processes governing the insurance and financial markets are assumed to be well-defined on this probability space.
@@ 2.1 Insurance Market
Consider a competitive insurance market that consists of a continuum of insurers, each offering insurance to individuals exposed to perfectly correlated risks. The idiosyncratic risks of individuals are ignored, as they can be diversified away. The cumulative loss process for a representative insurer, denoted by $L\triangleq\{L_{t}:t\geq 0\}$, evolves as the following dynamics:
@@
where $l>0$ represents the expected instantaneous loss rate, $\eta>0$ denotes the exposure level to the systematic loss risk, and $W^{I}\triangleq\{W^{I}_{t}:t\geq 0\}$ is a one-dimensional standard Brownian motion modeling the risk. [^11] In principle, the systematic loss risk may stem from multiple sources and can be modeled as multi-dimensional Brownian motion. For simplicity, and under the assumption of homogeneity in insurers’ risk exposures, we aggregate them into a single dimension for tractability.
@@
Specifically, on the supply side, insurers determine their underwriting amount based on the prevailing insurance price. The underwriting amount at time $t$, denoted by $x_{t}\geq 0$, constitutes the underwriting process $X\triangleq\{x_{t}:t\geq 0\}$. Then the insurer’s surplus process $M=\{m_{t}:t\geq 0\}$ evolves as the following dynamics:
@@
and the equilibrium price $\theta^{\ast}_{t}$ is solved endogenously. The price process is denoted by $\Theta\triangleq\{\theta_{t}:t\geq 0\}$, while the demand process is denoted by $D\triangleq\{d_{t}=d(\theta_{t}):t\geq 0\}$.
@@ 2.2 Financial Market
where $r$ is the risk-free rate, $\mu>r$ represents the expected return on the risky asset, and $\sigma>0$ denotes the volatility of the return. Here, $W^{S}\triangleq\{W^{S}_{t}:t\geq 0\}$ is a one-dimensional standard Brownian motion, which may be correlated with the insurance loss risk $W^{I}$. The correlation coefficient is given by $\rho\in(-1,1)$, such that $\mathrm{d}W^{I}_{t}\mathrm{d}W^{S}_{t}=\rho\mathrm{d}t$. Here we exclude the possibility of $\rho=\pm 1$ to simplify our analysis. In practice, perfect correlation between insurance and financial risks is almost impossible, as that would imply that one of the two markets would lose its distinct purpose. But we can still analyze the implications of such extreme cases from a theoretical perspective, see Subsection 3.7.
@@
Suppose that the insurer follows an investment strategy $Y\triangleq\{y_{t}:t\geq 0\}$, putting $y_{t}$ amount in the risky asset at time $t$, and the remaining $m_{t}-y_{t}$ in the risk-free asset. As a result, the insurer’s wealth process $M\triangleq\{m_{t}:t\geq 0\}$ follows:
$\mathrm{d}m_{t}=(x_{t}\theta_{t}l+y_{t}(\mu-r)+m_{t}r)\mathrm{d}t+x_{t}\eta\mathrm{d}W^{I}_{t}+y_{t}\sigma\mathrm{d}W^{S}_{t}.$ (2.1)